# VidLaDA: Bidirectional Diffusion Large Language Models for Efficient Video Understanding

### v1

@@ 4.2 Multi-modal Asynchronous Refreshing Strategy
Formally, a hidden state $H^{t,g}$ is updated only if $t\mod\tau_{(g,m)}=0$; otherwise, the cached state is reused. To preserve feature consistency across depths, we constrain the refresh interval of a shallow group to be a fixed integer multiple of the deeper group (i.e., $\tau_{(g-1,*)}=k\cdot\tau_{(g,*)}$). This enforces a synchronization constraint: whenever a stable shallow layer is updated, the volatile deep layers are simultaneously refreshed, ensuring that low-level features are always synchronized with high-level abstractions.
@@
Formally, we define $F_{n}$ as the subset of visual tokens belonging to the $n$-th frame. Accordingly, the local temporal neighborhood for frame $n$ is defined as $\mathcal{N}(F_{n})=F_{n-1}\cup F_{n}\cup F_{n+1}$. The attention for visual tokens within frame $n$ is then computed as:
$O_{[F_{n}]}=\text{Softmax}\left(\frac{Q_{[F_{n}]}K_{[\mathcal{N}(F_{n}),P]}^{\top}}{\sqrt{d_{k}}}\right)V_{[\mathcal{N}(F_{n}),P]},$ (2)
where $P$ denote the sets of text tokens of the prompt, $d_{k}$ represents the dimensionality of the key, $O_{[F_{n}]}$ represents the attention outputs for frame $n$, and the subscript $\mathcal{N}(F_{n}),P]$ denotes the concatenation of tokens from the neighboring frames and the text sequence. This operation effectively reduces the total complexity of the visual attention component from quadratic $\mathcal{O}(|V|^{2})$ to linear $\mathcal{O}(|\mathcal{N}(F_{n})|\times|V|)$.
@@
Adaptive Anchor Token Searching. However, naively restricting global interactions through chunking compromises the long-range information flow, leading to severe performance degradation (as analyzed in Appendix E.4.1). We hypothesize that specific anchor tokens within each frame act as critical inter-frame bridges, aggregating and transmitting global visual semantics across the temporal sequence. Exclusively relying on frame-wise chunk attention impedes these pathways, disrupting the hierarchical propagation of visual information, particularly in deeper layers. To restore global connectivity efficiently, we propose an optimized searching strategy comprising the following key components: (i) Adaptive Proxy Scoring: Computing the full attention map to identify anchor tokens is cost-prohibitive. Instead, we employ equidistant subsampling. Specifically, we sample a small subset of query tokens $\hat{S}$ (e.g., 32) to compute a low-rank proxy attention matrix:
$\hat{A}=\text{Softmax}(Q_{[\hat{S}]}K_{[V]}^{\top}/\sqrt{d_{k}})\in\mathbb{R}^{|\hat{S}|\times|V|}.$ (3)
To mitigate self-attention biases, we mask the diagonal entries where queries attend to themselves, yielding the debiased attention matrix $A$. (ii) Temporal Reuse: Leveraging Obs. 1, we perform this proxy scoring only at the first decoding step ($t=1$). The identified sink indices are cached and reused for all subsequent sparse attention steps, reducing the extra computational overhead. (iii) Group-wise Allocation: Leveraging the hierarchical consistency of anchor tokens in Obs. 1, we avoid per-layer searching. Instead, we search once per layer group. We allocate a decreasing budget $k_{g}$ for deeper groups (e.g., $k_{g-1}>k_{g}>\dots$). For group $g$, we aggregate the importance scores from the proxy matrix $A_{g_{\text{start}}}$ (computed at the group’s first layer) and select the top-$k_{g}$ indices.
To ensure a contiguous memory layout and consistent computational load, we enforce a fixed budget of anchor tokens per frame. For each frame $F_{j}$, we aggregate the importance scores across the sampled queries. Consequently, at step $t=1$, the anchor indices $\mathcal{I}_{(g,F_{j})}$ for each frame within group $g$ are determined by:
$A^{\prime}=\sum_{i}A_{g_{\text{start}}}[:,F_{j}]\in\mathbb{R}^{|F_{j}|},\,\mathcal{I}_{(g,F_{j})}=\text{Top}_{k_{g}}\left(A^{\prime}\right).$ (4)
The final set of global anchor tokens, $\mathcal{I}_{g}=\bigcup_{j}\mathcal{I}_{(g,F_{j})}$, is subsequently made visible to all tokens in the sequence, while these anchor tokens retain the ability to attend to all other tokens, effectively restoring global information propagation. Finally, exploiting the permutation invariance of bidirectional attention, we can relocate the anchor tokens within each frame to the beginning of the sequence to optimize memory access patterns.
@@ 5 Experiments
5.1 Experimental Setup
In this section, we first evaluate VidLaDA against state-of-the-art (SOTA) AR- and DLM-Based Video LLMs across diverse benchmarks. Then, we investigate the performance of CoT inference combined with MARS-Cache. Finally, this section concludes with a series of ablation studies. Detailed experimental setups are provided in Appendix F.1.
@@ 5.2 Main Results
We benchmark VidLaDA against current leading open-sourced AR-Based models and SOTA DLM baselines. The comparative results are presented in Table 1.
@@
w/o KV-Cache - 0.7 - 0.7 - 0.8
LLaDA-V (CoT) 63.4 3.5 42.5 4.3 55.7 4.5
+ Dual-Cache 64.2 23.8 43.2 28.3 57.1 27.2
+ MARS-Cache 63.4 35.2 43.6 40.0 57.7 37.9
VidLaDA (CoT) 67.4 2.7 49.0 3.3 59.8 2.1
@@ 5.3 Results under CoT Inference
To unlock the complex reasoning capabilities of models, we employ a structured Chain-of-Thought (CoT) inference pipeline inspired by (Zhou et al., 2024). This pipeline proceeds through four distinct stages: (1) task prompt routing, (2) reasoning analysis, (3) self-reflection, and (4) final answer generation. Further details on the specific prompt designs and experimental setup are provided in Appendix D.2 and Appendix F.4, respectively.
We evaluate the efficiency and robustness of our proposed strategies on EgoSchema (subset), MLVU (test), and LongVideoBench. Quantitative results are summarized in Table 2, with the batch size set to 1. To ensure practical inference speeds and establish a strong baseline, we employ parallel decoding (Wu et al., 2025) by default across all DLM settings. The primary objective of this experiment is to verify the generalizability of MARS-Cache on DLMs (LLaDA-V and VidLaDA) under CoT inference, and to benchmark its efficiency against the AR baseline.
@@
Overall, these results demonstrate that MARS-Cache substantially accelerates CoT-style video reasoning, generalizes well across DLM-Based Video LLMs, and achieves competitive (or even better) efficiency than the strong AR baseline while maintaining comparable reasoning performance.
@@ 5.4 Ablation Studies
In this section, we summarize the key conclusions from our ablation studies on the MARS-Cache design. Comprehensive experimental results are provided in Appendix E.4. (i) The Necessity of Anchor Token: Our assessment of global connectivity reveals that retaining anchor tokens is crucial; completely reverting to frame-wise chunk attention significantly degrades performance. We find that preserving full attention in shallow layers to maintain structural integrity while aggressively pruning anchors in deep layers achieves the optimal trade-off between reasoning accuracy and computational cost. (ii) Asynchronous Refreshing Strategy: Leveraging the observation that visual hidden states exhibit higher temporal stability than textual ones, we implement a modality-wise refreshing strategy. We conclude that setting the visual refresh interval larger than the textual interval (optimally $R_{v/t}\approx 2$) significantly boosts throughput without compromising accuracy, whereas excessive ratios lead to degradation. Additionally, regarding network depth, a pyramid-style schedule (increasing update frequency for deeper, more volatile layers) consistently outperforms uniform schedules and is orthogonal to the modality-wise refreshing strategy. (iii) Search Overhead: For the adaptive anchor searching mechanism, we determine that a query subset size of 128 offers a robust compromise. It is sufficient to identify high-quality global information hubs while keeping the computational overhead negligible, whereas using the full sequence leads to memory bottlenecks.
@@ 6 Conclusion
Appendix A Detailed Problem Formulation and Background
@@
A.1 Autoregressive-Based Video Language Models
The prevailing paradigm for MLLMs treats video understanding as a conditional sequence generation task. Given visual tokens $V$ (derived from video $\mathcal{V}$ via a ViT and projector) and prompt tokens $P$, the model maximizes the conditional likelihood of the response $R$ autoregressively:
$\mathcal{L}_{\text{AR}}=-\sum_{i=1}^{|R|}\log p_{\theta}(r_{i}\mid V,P,R_{<i}),$ (5)
where $p_{\theta}$ is parameterized by a Transformer Decoder (Vaswani et al., 2017; Grattafiori et al., 2024). This approach necessitates a causal attention mask $M$, defined as:
$M_{ij}=\begin{cases}0&\text{if }j\leq i,\\ -\infty&\text{otherwise.}\end{cases}$ (6)
@@ A.1 Autoregressive-Based Video Language Models
A.2 Diffusion Language Models (DLMs)
@@
Forward Process. The forward process adds noise by independently masking tokens. Let $x_{0}$ denote the clean sequence (containing conditions $V,P$, and response $R$). At continuous time $t\in[0,1]$, each token $x_{t}^{(i)}$ in the response is replaced by [MASK] with probability $t$. The transition distribution is:
$q_{t|0}(x_{t}\mid x_{0})=\prod_{i}q_{t|0}(x_{t}^{(i)}\mid x_{0}^{(i)}),$ (7)
$\text{where }q_{t|0}(x_{t}^{(i)}\mid x_{0}^{(i)})=\begin{cases}1-t&\text{if }x_{t}^{(i)}=x_{0}^{(i)},\\ t&\text{if }x_{t}^{(i)}=\texttt{[MASK]}.\end{cases}$ (8)
Reverse Process and Objective. The reverse process aims to recover $x_{0}$ from the masked state $x_{t}$. A neural network $p_{\theta}(\cdot\mid x_{t})$ predicts the original tokens for masked positions. In visual instruction tuning, we condition on unmasked $V$ and $P$, applying masking only to $R$. The objective minimizes the variational upper bound:
$\mathcal{L}_{\text{DLM}}=-\mathbb{E}_{t,R_{0},R_{t}}\left[\frac{1}{t}\sum_{i=1}^{|R|}m_{t}^{(i)}\log p_{\theta}(R_{0}^{(i)}\mid V,P,R_{t})\right],$ (9)
where $m_{t}^{(i)}=\mathbb{I}[R_{t}^{(i)}=\texttt{[MASK]}]$ is the mask indicator.
Bidirectional Attention. Unlike AR models, the prediction of $R_{0}^{(i)}$ and understanding of $V,P,R$ in DLMs depend on the global context. Thus, we utilize full bidirectional attention:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V.$ (10)
This mechanism enables global spatiotemporal modeling for $V$. During inference, the model starts from a fully masked $R_{1}$ and iteratively refines the sequence by predicting $R_{0}$ and re-masking based on a schedule.
Appendix B Related Work
B.1 Autoregressive-Based Video Large Language Models
Following the success of Large Language Models (LLMs) in text generation, Multimodal Large Language Models (MLLMs) have rapidly evolved to perceive visual inputs. Early works like LLaVA (Liu et al., 2024b) projected static visual features into the LLM’s token space, enabling image-text question answering. This paradigm was quickly adapted to the video domain, deriving the Video Large Language Models (Video LLMs). Models such as Video-LLaVA (Lin et al., 2024), VideoChat (Li et al., 2025a), VideoLLaMA (Cheng et al., 2024), LongVILA (Chen et al., 2024c), InternVideo (Wang et al., 2024b), mPLUG-Owl (Ye et al., 2024), and PLLaVA (Xu et al., 2024) aggregate temporal frames using pooling or Q-Former (Li et al., 2023a) structures before feeding them into an autoregressive LLM decoder. More recent state-of-the-art models, including LLaVA-OneVision (Li et al., 2024a), LLaVA-Video (Zhang et al., 2024b), Qwen2-VL (Wang et al., 2024a), and Qwen2.5-VL (Bai et al., 2025b) scale up the visual resolution and context length, achieving impressive performance on general video benchmarks.

### v2

@@ 4.2 Multi-modal Asynchronous Refreshing Strategy
Formally, a hidden state ${\bm{H}}^{t,g}$ is updated only if $t\mod\tau_{(g,m)}=0$; otherwise, the cached state is reused. To preserve feature consistency across depths, we constrain the refresh interval of a shallow group to be a fixed integer multiple of the deeper group (i.e., $\tau_{(g-1,*)}=k\cdot\tau_{(g,*)}$). This enforces a synchronization constraint: whenever a stable shallow layer is updated, the volatile deep layers are simultaneously refreshed, ensuring that low-level features are always synchronized with high-level abstractions.
@@
Formally, we define $\mathcal{F}_{n}$ as the subset of visual tokens ${\bm{E}}^{\mathcal{V}}$} belonging to the $n$-th frame. Accordingly, the local temporal neighborhood for frame $n$ is defined as $\mathcal{N}(\mathcal{F}_{n})=\mathcal{F}_{n-1}\cup\mathcal{F}_{n}\cup\mathcal{F}_{n+1}$. The attention for visual tokens within frame $n$ is then computed as:
${\bm{O}}_{[\mathcal{F}_{n}]}=\text{Softmax}\left(\frac{{\bm{Q}}_{[\mathcal{F}_{n}]}{\bm{K}}_{[\mathcal{N}(\mathcal{F}_{n})]}^{\top}}{\sqrt{d_{k}}}\right){\bm{V}}_{[\mathcal{N}(\mathcal{F}_{n})]},$ (2)
where $d_{k}$ represents the dimensionality of the key, and the ${\bm{O}}_{[\mathcal{F}_{n}]}$ represents the attention outputs for frame $n$. This operation effectively reduces the total complexity of the attention mechanism between video tokens from quadratic $\mathcal{O}(|{{\bm{E}}^{\mathcal{V}}}|^{2})$ to linear $\mathcal{O}(|\mathcal{N}(\mathcal{F}_{n})|\times|{{\bm{E}}^{\mathcal{V}}}|)$.
@@
Adaptive Anchor Token Searching. However, naively restricting global interactions through chunking compromises the long-range information flow, leading to severe performance degradation (as analyzed in Appendix F.5.1). We hypothesize that specific anchor tokens within each frame act as critical inter-frame bridges, aggregating and transmitting global visual semantics across the temporal sequence. Exclusively relying on frame-wise chunk attention impedes these pathways, disrupting the hierarchical propagation of visual information, particularly in deeper layers. To restore global connectivity efficiently, we propose an optimized searching strategy comprising the following key components: (i) Adaptive Proxy Scoring: Computing the full attention map to identify anchor tokens is cost-prohibitive. Instead, we employ equidistant subsampling. Specifically, we sample a small subset of query tokens $\hat{\mathcal{S}}$ (e.g., $|\hat{\mathcal{S}}|=32$) to compute a low-rank proxy attention matrix:
$\hat{{\bm{A}}}=\text{Softmax}({\bm{Q}}_{[\hat{\mathcal{S}}]}{\bm{K}}_{[\mathcal{S^{\mathcal{V}}}]}^{\top}/\sqrt{d_{k}})\in\mathbb{R}^{|\hat{\mathcal{S}}|\times|\mathcal{S^{\mathcal{V}}}|},$ (3)
where the $\mathcal{S^{\mathcal{V}}}$ denotes the set of video tokens. To mitigate self-attention bias, we mask the entries where queries attend to themselves, yielding the debiased attention matrix $\bm{A}$}. (ii) Temporal Reuse: Leveraging Obs. 1, we perform this proxy scoring only at the first decoding step ($t=1$). The identified sink indices are cached and reused for all subsequent sparse attention steps, reducing the extra computational overhead. (iii) Group-wise Allocation: Leveraging the hierarchical consistency of anchor tokens in Obs. 1, we avoid per-layer searching. Instead, we search once per layer group. We allocate a decreasing budget $k_{g}$ for deeper groups (e.g., $k_{g-1}>k_{g}>\dots$). For group $g$, we aggregate the importance scores from the proxy matrix ${\bm{A}}^{g_{\text{start}}}$ (computed at the group’s first layer) and select the top-$k_{g}$ indices.
To ensure a contiguous memory layout and consistent computational load, we enforce a fixed budget of anchor tokens per frame. For each frame $\mathcal{F}_{j}$, we aggregate the importance scores across the sampled queries. Consequently, at step $t=1$, the anchor indices $\mathcal{I}_{(g,\mathcal{F}_{j})}$ for each frame within group $g$ are determined by:
${\bm{A}}^{\prime}=\sum_{i}{\bm{A}}^{g_{\text{start}}}_{[i,\mathcal{F}_{j}]}\in\mathbb{R}^{|\mathcal{F}_{j}|},\,\mathcal{I}_{(g,\mathcal{F}_{j})}=\text{Top}_{k_{g}}\left({\bm{A}}^{\prime}\right).$ (4)
The final set of global anchor tokens, $\mathcal{I}_{g}=\bigcup_{j}\mathcal{I}_{(g,\mathcal{F}_{j})}$, is subsequently made visible to all tokens in the sequence, while these anchor tokens retain the ability to attend to all other tokens, effectively restoring global information propagation. Finally, exploiting the permutation invariance of bidirectional attention, we can relocate the anchor tokens within each frame to the beginning of the sequence to optimize memory access patterns.
@@ 5 Experiments
In this section, we first evaluate VidLaDA against state-of-the-art (SOTA) AR and DLM baselines across diverse benchmarks. Then, we investigate the performance of Chain-of-Thought (CoT) inference combined with MARS-Cache. Finally, this section concludes with a series of ablation studies. Detailed experimental setups are provided in Appendix G.1.
@@ 5.1 Main Results
We benchmark VidLaDA against the current leading open-source SOTA AR and DLM baselines. The comparative results are presented in Table 1.
@@
LLaDA-V (CoT) 65.4 3.5 43.2 4.3 58.1 4.5
+ Dual-Cache 64.4 23.8 43.2 28.3 59.2 27.2
+ MARS-Cache 65.2 35.2 43.6 40.0 59.6 37.9
VidLaDA (CoT) 67.4 2.7 50.2 3.3 59.8 2.1
@@ 5.2 Results under CoT Inference
To unlock the complex reasoning capabilities of models, we employ a structured Chain-of-Thought (CoT) inference pipeline inspired by (Zhou et al., 2024). This pipeline proceeds through four distinct stages: (1) task prompt routing, (2) reasoning analysis, (3) self-reflection, and (4) final answer generation. Further details on the specific prompt designs and experimental setup are provided in Appendix E.2 and Appendix G.4, respectively.
We evaluate the efficiency and robustness of our proposed strategies on EgoSchema (subset), MLVU (test), and LongVideoBench. Quantitative results are summarized in Table 2, with the batch size set to 1. To ensure practical inference speeds and establish a strong baseline, we employ parallel decoding (Wu et al., 2025) by default across all DLM settings. The primary objective of this experiment is to verify the generalizability of MARS-Cache on DLMs (LLaDA-V and VidLaDA) under CoT-style inference, and to benchmark its efficiency against the AR baseline.
@@
Overall, these results demonstrate that MARS-Cache substantially accelerates CoT-style video reasoning and generalizes well across DLM-based Video LLMs. Some visual examples of the reasoning trajectories generated with MARS-Cache are detailed in Appendix F.3.
@@ 5.3 Ablation Studies
In this section, we summarize the key conclusions from our ablation studies. Comprehensive results about MARS-Cache (iiiv) are in Appendix F.5. (i) Architectures. To further rule out training confounders, we compare AR and DLM baselines under matched fine-tuning settings (Appendix F.4). We find that, given identical data, DLM outperforms AR, lending support to an architectural advantage. (ii) The Necessity of Anchor Token: Our assessment of global connectivity reveals that retaining anchor tokens is crucial; completely reverting to frame-wise chunk attention significantly degrades performance. We find that preserving full attention in shallow layers to maintain structural integrity while aggressively pruning anchors in deep layers achieves the optimal trade-off between reasoning accuracy and computational cost. (iii) Asynchronous Refreshing Strategy: Leveraging the observation that visual hidden states exhibit higher temporal stability than textual ones, we implement a modality-wise refreshing strategy. We conclude that setting the visual refresh interval larger than the textual interval (optimally $R_{v/t}\approx 2$) significantly boosts throughput without compromising accuracy, whereas excessive ratios lead to degradation. Additionally, regarding network depth, a pyramid-style schedule (increasing update frequency for deeper, more volatile layers) consistently outperforms uniform schedules and is orthogonal to the modality-wise refreshing strategy. (iv) Search Overhead: For the adaptive anchor searching mechanism, we determine that a query subset size of 128 offers a robust compromise. It is sufficient to identify high-quality global information hubs while keeping the computational overhead negligible, whereas using the full sequence leads to memory bottlenecks.
@@ 6 Conclusion
Appendix A Main Notation Introduction
Table 3: Meanings of the Main Notations
Notation Size Meaning
|||*| Constant The length of the corresponding sequence.
$d$ Constant The model dimension.
$d_{k}$ Constant The dimensionality of the key in the multi-head mechanism.
$\mathcal{V}$} $F\times C\times H\times W$ The input video frames.
${\bm{E}}^{\mathcal{V}}$} $|{{\bm{E}}^{\mathcal{V}}}|\times d$ The spatiotemporal visual tokens encoded by the ViT.
$P$ $|P|$ The token IDs of the text prompt produced by the tokenizer.
$R$ $|R|$ The token IDs of the text response produced by the tokenizer.
${\bm{z}}_{t}$ $d$ The latent representation encoded by ViT and LLM at timestep $t$.
$\mathcal{I}$ $|\mathcal{I}|$ The index set of anchor tokens.
$\tau_{(g,m)}$ Constant The refresh interval in group $g$ for modality $m$.
$\bm{H}$} $|{\bm{H}}|\times d$ The hidden states in the LLMs.
$\mathcal{F}_{n}$ $|\mathcal{F}_{n}|$ The index set of visual tokens ${\bm{E}}^{\mathcal{V}}$} belonging to the $n$-th frame.
$\bm{O}$} $|{\bm{O}}|\times d_{k}$ The attention output matrix.
$\bm{Q}$} $|{\bm{Q}}|\times d_{k}$ The attention query matrix.
$\bm{K}$} $|{\bm{K}}|\times d_{k}$ The attention key matrix.
$\bm{V}$} $|{\bm{V}}|\times d_{k}$ The attention value matrix.
[…]
$\mathcal{S^{\mathcal{V}}}$ $|\mathcal{S^{\mathcal{V}}}|$ The index set of video tokens.
Appendix B Detailed Problem Formulation and Background
@@
B.1 Autoregressive-Based Video Language Models
The prevailing paradigm for Multimodal Large Language Models (MLLMs) treats video understanding as a conditional sequence generation task. Given visual tokens ${\bm{E}}^{\mathcal{V}}$} (derived from video $\mathcal{V}$} via a ViT and projector) and prompt tokens $P$, the model maximizes the conditional likelihood of the response $R$ autoregressively:
$\mathcal{L}_{\text{AR}}=-\sum_{i=1}^{|R|}\log p_{\theta}(R_{i}\mid{{\bm{E}}^{\mathcal{V}}},P,R_{<i}),$ (5)
where $p_{\theta}$ is parameterized by a Transformer Decoder (Vaswani et al., 2017; Grattafiori et al., 2024). This approach necessitates a causal attention mask $\bm{M}$}, defined as:
${\bm{M}}_{ij}=\begin{cases}0&\text{if }j\leq i,\\ -\infty&\text{otherwise.}\end{cases}$ (6)
@@ B.1 Autoregressive-Based Video Language Models
B.2 Diffusion Language Models (DLMs)
@@
Forward Process. In the context of conditional generation, the forward process adds noise to the response $R$ while keeping the conditions (visual inputs $\mathcal{V}$} and text prompt $P$) unmasked. Let $R^{0}$ denote the clean response sequence. At continuous time $t\in[0,1]$, each token $R_{i}^{t}$ in the response is independently replaced by [MASK] with probability $t$. The transition distribution is defined specifically over the response tokens:
$q^{t|0}(R^{t}\mid R^{0})=\prod_{i=1}^{|R|}q^{t|0}(R_{i}^{t}\mid R_{i}^{0}),$ (7)
$\text{where }q^{t|0}(R_{i}^{t}\mid R_{i}^{0})=\begin{cases}1-t&\text{if }R_{i}^{t}=R_{i}^{0},\\ t&\text{if }R_{i}^{t}=\texttt{[MASK]}.\end{cases}$ (8)
Reverse Process and Objective. The reverse process aims to recover the clean response $R^{0}$ from the masked state $R^{t}$. A neural network $p_{\theta}(\cdot\mid{{\bm{E}}^{\mathcal{V}}},P,R^{t})$ predicts the original tokens for masked positions within the response, conditioned on the fully observable ${\bm{E}}^{\mathcal{V}}$} and $P$. The objective minimizes the variational upper bound:
$\mathcal{L}_{\text{DLM}}=-\mathbb{E}_{t,R^{0},R^{t}}\left[\frac{1}{t}\sum_{i=1}^{|R|}m_{i}^{t}\log p_{\theta}(R_{i}^{0}\mid{{\bm{E}}^{\mathcal{V}}},P,R^{t})\right],$ (9)
where $m_{i}^{t}=\mathbb{I}[R_{i}^{t}=\texttt{[MASK]}]$ is the mask indicator.
Bidirectional Attention. Unlike AR models, the prediction of $R_{i}^{0}$ and understanding of ${{\bm{E}}^{\mathcal{V}}},P,R$ in DLMs depend on the global context. Thus, we utilize full bidirectional attention:
$\text{Attention}({\bm{Q}},{\bm{K}},{\bm{V}})=\text{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\right){\bm{V}}.$ (10)
This mechanism enables global spatiotemporal modeling for ${\bm{E}}^{\mathcal{V}}$}. During inference, the model starts from a fully masked $R^{1}$ and iteratively refines the sequence by predicting $R^{0}$ and re-masking based on a schedule.
Appendix C Related Work
C.1 Autoregressive-Based Video Large Language Models
Following the success of Large Language Models (LLMs) in text generation, Multimodal Large Language Models (MLLMs) have rapidly evolved to perceive visual inputs. Early works like LLaVA (Liu et al., 2024b) projected static visual features into the LLM’s token space, enabling image-text question answering. This paradigm was quickly adapted to the video domain, deriving the Video Large Language Models (Video LLMs). Models such as Video-LLaVA (Lin et al., 2024), VideoChat (Li et al., 2025a), VideoLLaMA (Cheng et al., 2024), LongVILA (Chen et al., 2024c), InternVideo (Wang et al., 2024c), mPLUG-Owl (Ye et al., 2024), VideoLoom (Shi et al., 2026), and PLLaVA (Xu et al., 2024) aggregate temporal frames using pooling or Q-Former (Li et al., 2023a) structures before feeding them into an autoregressive LLM decoder. More recent state-of-the-art models, including LLaVA-OneVision (Li et al., 2024a), LLaVA-Video (Zhang et al., 2024b), ProLongVid (Wang et al., 2025b), Qwen2-VL (Wang et al., 2024a), and Qwen2.5-VL (Bai et al., 2025b) scale up the visual resolution and context length, achieving impressive performance on general video benchmarks.